OpenStudy (societyyay):

Guys how do I prove this identity ---> 4+sinfeta = 6(cos square)feta

OpenStudy (societyyay):

Sorry solve the equation of this where 0 ≤0 ≤ 360

OpenStudy (mrnood):

theta

OpenStudy (societyyay):

Yeah theta**

OpenStudy (3mar):

I think that is how your equation looks like! \[\Large 4+\sin \theta=6\cos^2\theta\] Am I right?

OpenStudy (societyyay):

Yess!

OpenStudy (3mar):

That is great! Any ideas how to start?

OpenStudy (societyyay):

Nope :/ none of the identities rule work at least

OpenStudy (3mar):

try this one! \[\Huge\color{Coral}{sin^2 \theta +cos^2\theta=1}\] or \[\Huge\color{MediumVioletRed }{6~sin^2 \theta +6~cos^2\theta=6}\]

OpenStudy (3mar):

substitute with \(6~cos^2\theta\) into your equation!

OpenStudy (societyyay):

I don't get it. How does that apply to the question?

OpenStudy (societyyay):

What about the 4 then?

OpenStudy (3mar):

I will show you!

OpenStudy (3mar):

 Did you get the idea?